# Revisiting Optimal Allocations for Binary Responses: Insights from Considering Type-I Error Rate Control

### v3

@@ 4.3 Simulation Results
Building on the theoretical results, we now analyze the simulation for the new allocation proportions. Figure 1 shows that the score test generally yields a lower type-I error rate than the Wald test, as seen in the comparison of CR for both tests ($\rho_{CR_{0}}$ vs. $\rho_{CR_{1}}$). The two optimal optimal proportions $\rho^{n}_{N_{0}}$ and $\rho^{n}_{R_{0}}$, successfully control the type-I error rate and even outperform CR in this regard.
Table 4 shows that $\rho^{n}_{N_{0}}$ and $\rho^{n}_{R_{0}}$ achieve similar power to CR, with differences within $2\%$ in either direction across parameter combinations. However, the difference in type-I error rate control is substantial: CR fails to control the nominal level, while both new proportions do. This does demonstrate that neither of the two designs using the new allocation proportion requires any of the potential solutions discussed in Section 3, including larger burn-in sizes. While $\rho^{n}_{N_{0}}$ may reduce the ENS compared to CR, $\rho^{n}_{R_{0}}$ tends to increase it and typically allocates more patients to the superior arm.
@@ 5.1 Early Phase Trial
Results are summarized in Table 5. Among the allocation proportions examined, only the RSHIR-like allocation, $\rho^{n}_{R_{0}}$, successfully controlled the type-I error rate. Even complete randomization exhibited inflation of type-I error. The Neyman allocation, $\rho^{n}_{N_{0}}$, for the score test reduces type-I error rate inflation compared to complete randomization, though some inflation remains. However, the Neyman and RSHIR methods resulted in highly inflated type-I error rates, making them impractical for use.
Our findings suggest that the RSHIR-like allocation is the only viable method, as it properly controls the type-I error rate. While it has slightly lower power than complete randomization, it provides a significant increase in the expected number of successes, making it a preferable approach in this context.
@@ 5.2 Confirmatory Trial: CALISTO
The trial results revealed a success rate of 94.1% in the placebo group and 99.1% in the treatment group. Originally, the study was designed to include 3002 patients, but only $n=1502$ patients were included in the analysis. We compare optimal proportions aimed at minimizing failures $\rho_{R_{1}}$ and $\rho^{n}_{R_{0}}$, as well as maximizing power, $\rho_{N_{1}}$ and $\rho^{n}_{N_{0}}$ to CR. The analysis involved comparing test statistics $Z_{0}$ and $Z_{1}$ for each method. For each method, a burn-in period of $2$ patients per arm was implemented.
Table 5 demonstrates the proportions $\rho_{CR}$, $\rho^{n}_{R_{0}}$ and $\rho^{n}_{N_{0}}$ effectively control the type-I error rate. While all designs have similar power, both proportions $\rho^{n}_{R_{0}}$ and $\rho^{n}_{N_{0}}$ are able to increase patient benefit significantly. Especially, using $\rho^{n}_{R_{0}}$ leads to 25 additional success compared to CR.
@@ 6 Discussion and Future Research
In this paper, we introduce two optimal allocation proportions that demonstrate strong type-I error rate control in adaptive two-armed trials with binary outcomes. In finite samples they achieve type-I error rate control for the largest margin of the parametric space (90%) and offer the best type-I error rate control for the remaining 10%. These proportions offer a promising balance between statistical rigor and practical relevance. Our proposed methods demonstrate power performance that is comparable to existing approaches, with occasional gains, particularly in small sample settings. Importantly, the proposed proportions have the advantage of increasing the likelihood that patients receive the superior treatment.
@@ Acknowledgments
The authors acknowledge funding and support from the UK Medical Research Council (grants MC UU 00002/15 and MC UU 00040/03 (SSV), as well as an MRC Biostatistics Unit Core Studentship (LP) and the Cusanuswerk e.V. (LP). SSV is part of PhaseV’s advisory board. Additionally, LP’s visit to the Department of Statistics at George Mason University to collaborate with WR on this project was supported by the Trials Methodology Research Partnership Doctoral Training Program and the Fisher Memorial Trust. LP is thankful for the departments hospitality.

### v4

@@ 4.3 Simulation Results
Building on the theoretical results, we now analyze the simulation for the new allocation proportions. Figure 1 shows that the score test generally yields a lower type-I error rate than the Wald test, as seen in the comparison of CR for both tests ($\rho_{CR_{0}}$ vs. $\rho_{CR_{1}}$). The optimal optimal proportion $\rho^{n}_{N_{0}}$ successfully controls the type-I error rate and even outperform CR in this regard while the optimal proportion $\rho^{n}_{R_{0}}$ shows similar (or even better) performance for majority of the parametric space but is slightly inflated for values around $0.8$.
Table 4 shows that $\rho^{n}_{N_{0}}$ and $\rho^{n}_{R_{0}}$ achieve similar power to CR, with differences within few percent in either direction across parameter combinations. The difference in type-I error rate control varies. The Neyman-like optimal allocation $\rho^{n}_{N_{0}}$ generally provides better control of the type-I error rate compared to CR, except for values near the boundaries (i.e., close to 0 or 1). The RSHIR-like allocation proportion $\rho^{n}_{R_{0}}$ shows even better type-I error rate control for the majority of the parametric space, but the type-I error rate is inflated earlier for values approaching 1. While $\rho^{n}_{N_{0}}$ may reduce the ENS compared to CR, $\rho^{n}_{R_{0}}$ tends to increase it and typically allocates more patients to the superior arm.
@@ 5.1 Early Phase Trial
Results are summarized in Table 5. Among the allocation proportions examined, the RSHIR-like allocation $\rho^{n}_{R_{0}}$, the Neyman-like allocation $\rho^{n}_{N_{0}}$, and complete randomization, when testing with $Z_{0}$, successfully control the type-I error rate. Complete randomization when testing with the Wald test exhibited slight inflation of type-I error. However, the original Neyman $\rho_{N_{1}}$ and RSHIR $\rho_{R_{1}}$ allocations resulted in highly inflated type-I error rates, making them impractical for use.
Our findings suggest that both the RSHIR-like and Neyman-like allocation are viable methods, as they properly control the type-I error rate. While they have slightly lower power (less thatn $1\%$ difference) than complete randomization, they provide a significant increase in the expected number of successes, making them preferable approaches in this context.
@@ 5.2 Confirmatory Trial: CALISTO
The trial results revealed a success rate of 94.1% in the placebo group and 99.1% in the treatment group. Originally, the study was designed to include 3002 patients, but only $n=1502$ patients were included in the analysis. We compare optimal proportions aimed at minimizing failures $\rho_{R_{1}}$ and $\rho^{n}_{R_{0}}$, as well as maximizing power, $\rho_{N_{1}}$ and $\rho^{n}_{N_{0}}$ to CR. For each method we use the test statistic $Z_{0}$ or $Z_{1}$ that was used in the optimization problem but for CR we present both. For each method, a burn-in period of $2$ patients per arm was implemented.
Table 5 demonstrates that CR does lead to asymptotic type-I error control using either the Wald or score test. The proportions $\rho^{n}_{R_{0}}$ and $\rho^{n}_{N_{0}}$ almost control type-I error rate, while $\rho_{R_{1}}$ and more so $\rho_{N_{1}}$ do not. While all designs have similar power, both proportions $\rho^{n}_{R_{0}}$ and $\rho^{n}_{N_{0}}$ are able to increase patient benefit significantly. Especially, using $\rho^{n}_{R_{0}}$ leads to 25 additional success compared to CR.
@@ 6 Discussion and Future Research
In this paper, we introduce two optimal allocation proportions that demonstrate type-I error rate control in adaptive two-armed trials with binary outcomes. In finite samples they achieve type-I error rate control for the largest margin of the parametric space (90%). These proportions offer a promising balance between statistical rigor and practical relevance. Our proposed methods demonstrate power performance that is comparable to existing approaches, with occasional gains, particularly in small sample settings. Importantly, the proposed proportions have the advantage of increasing the likelihood that patients receive the superior treatment.
@@ Acknowledgments
The authors acknowledge funding and support from the UK Medical Research Council (grants MC UU 00002/15 and MC UU 00040/03 (SSV), as well as an MRC Biostatistics Unit Core Studentship (LP) and the Cusanuswerk e.V. (LP). SSV is part of PhaseV’s advisory board. Additionally, LP’s visit to the Department of Statistics at George Mason University to collaborate with WR on this project was supported by the Trials Methodology Research Partnership Doctoral Training Program and the Fisher Memorial Trust. LP is thankful for the departments hospitality. We thank Dr. Stef Baas for revising our codes and spotting two minor errors. The authors are grateful to the AE and referees for their insightful suggestions.